OpenStudy (anonymous):

find cos theta if theta is an angle in standard position and the point with coordinates (3,-4) lies on the terminal side of the angle

OpenStudy (anonymous):

 is that right to start off?

OpenStudy (anonymous):

yes but the cos angle

OpenStudy (anonymous):

cos theta= x/c or cot theta =x/y

OpenStudy (anonymous):

okay so is the angle in the wrong place or